OpenStudy (anonymous):

solve equations by elimination -2x-2y=3 x-2y=1

OpenStudy (texaschic101):

-2x - 2y = 3 x - 2y = 1 --->(2) --------------------- -2x - 2y = 3 2x - 4y = 2 (result of multiplying by 2) --------------------add -6y = 5 y = -5/6 now sub -5/6 in for y in either equation to find x x - 2y = 1 x - 2(-5/6) = 1 x + 10/6 = 1 x + 5/3 = 1 x = 1 - 5/3 x = 3/3 - 5/3 x = -2/3 check... -2x - 2y = 3 -2(-2/3) - 2(-5/6) = 3 4/3 + 10/6 = 3 8/6 + 10/6 = 3 18/6 = 3 3 = 3 (correct) x = -2/3 and y = -5/6